# Right Reward Right Time for Federated Learning

### v1

@@ Abstract
Critical learning periods (CLPs) in federated learning (FL) refer to early stages during which low-quality contributions (e.g., sparse training data availability) can permanently impair the learning performance of the global model owned by the model owner (i.e., the cloud server). However, strategies to motivate clients with high-quality contributions to join the FL model training process and share trained model updates during CLPs remain under-explored. Additionally, existing incentive mechanisms in FL treat all training periods equally, which consequently fails to motivate clients to participate early. Compounding this challenge is the cloud’s limited knowledge of client training capabilities due to privacy regulations, leading to information asymmetry. Thus, in this article, we propose a time-aware incentive mechanism, called Right Reward Right Time (R3T), to encourage client involvement, especially during CLPs, to maximize the utility of the cloud server in FL. Specifically, the cloud utility function captures the trade-off between the achieved model performance and payments allocated for clients’ contributions, while accounting for the client’s time and system capabilities, the efforts, joining time, and rewards. Then, we analytically derive the optimal contract for the cloud and devise a CLPs-aware mechanism to incentivize early participation and efforts while maximizing cloud utility, even under information asymmetry. By providing the right reward at the right time, our approach can attract the highest-quality contributions during CLPs. Simulation and proof-of-concept studies show that R3T increases cloud utility and is more economically effective than benchmarks. Notably, our proof-of-concept results demonstrate up to a 47.6% reduction in the total number of clients and up to a 300% improvement in convergence time while reaching competitive test accuracies compared with incentive mechanism benchmarks.
Index Terms:
Blockchain, Contract theory, Critical learning periods, Federated learning, Incentive mechanism.
1 Introduction
1.1 Background and Motivations
@@
Federated learning (FL) is a privacy-preserving data sharing and learning paradigm, where distributed clients (e.g., devices or organizations) collaboratively train an artificial intelligence model without disclosing their private data [1, 2]. In FL settings, the importance of all training rounds and the efforts from distributed clients across these rounds are regarded as equally important. However, this setting is questioned by a recent discovery of a phenomenon in FL known as critical learning periods (CLPs), which highlight the varying impacts of different training rounds and significantly affect FL learning and training efficiency [3]. This finding was inspired by earlier research on CLPs in species’ cognitive and learning functions (i.e., humans and animals) and deep learning networks within a centralized setting [4, 5, 6, 7]. Particularly, CLPs are early learning phases (c.f., Fig. 1(1(a))) profoundly affect long-term learning capabilities [8]. Empirical studies show that the absence of essential and proper learning experiences during these periods can lead to irreversible deficits. For instance, barn owls exposed to misaligned auditory and visual cues during their critical development periods fail to properly localize spatial locations in adulthood [9]. This phenomenon was also observed in deep neural network, in which any deficits and low effort (e.g., sparse data availability, late client participation) during this initial learning stage can cause a lasting impaired effect on learning model performance, regardless of subsequent efforts (c.f., Fig. 1(1(b))). The reason for the phenomenon is still unclear, but possible explanations include the depth of the learning model, structure of training data distribution, or defects in the implementation and training process [5, 6, 7].
@@
Figure 1: Illustration of: a) CLPs in the biological field [10], and b) how early effort affects long-term FL model performance.
@@
Despite the promising merits of enhancing FL learning and training efficiency (e.g., accuracy and convergence rate), CLPs in FL remain under-explored, including a lack of transparent, fair, and timely incentives for participation and contribution enticement and information asymmetry among clients and the global model owner (e.g., cloud server) in these periods. Current existing CLP-considered works in FL, such as CLPs detection, client selection, or CLP-aware robustness, are based on a common assumption of voluntary engagement of clients in the training process (e.g., [3, 11, 12, 13]). However, this is impractical because clients are rational, strategic, and self-interested and need sufficient rewards to compensate for their efforts (e.g., data, computation resources, and time costs). This leads to a lack of client motivation for contribution, resulting in the unsustainable development of FL systems. Although several works have been proposed by authors (e.g., [14, 15, 16, 17]) to bring incentivization by considering effort-reward approaches to fairly compensate FL clients’ effort, they treat clients’ efforts and their impacts on the global model performance on every training round equally and ignore the temporal factor, whereas efforts in CLPs have more weights and mainly contribute to the overall performance of the FL global model. By failing to prioritize CLP-driven incentive solutions, current methods inefficiently allocate cloud budgets (e.g., monetary rewards), thereby reducing cloud utility—defined as the difference between global model performance and compensation costs (which is defined later in Section 2).
Additionally, there are significant issues of information asymmetry between FL clients and the cloud server, which make CLPs misaligned and cause disputes among the cloud server [^1] The cloud server and the cloud are used interchangeably. and clients. Specifically, due to privacy restrictions, the cloud server lacks prior knowledge of the clients’ training capabilities, including the computational resources, data properties (e.g., data quantity or distribution), or available joining time. Therefore, assumptions that the cloud server knows precise information [18, 19, 20], such as clients’ computing resources or clients’ data size, may not hold in practice. This uncertainty can adversely affect the alignment of CLPs in determining the final model performance in FL. For example, under conditions of information asymmetry, the cloud server randomly selects clients with different training capacities, resulting in the underrepresentation of high-quality clients [^2] We refer high-quality clients to clients with high-quality contributions and efforts, such as a huge amount of training data and large computation resources, and they are available in CLPs., thereby high-quality contributions and efforts may get diluted with low-quality ones when averaged together, impairing the early advantage impact of CLPs on the final model performance. There are also uncertainty regarding whether clients will exert the promised effort post-contract [21]. Despite the presence of contractually agreed-upon incentives, FL clients may still engage in dishonest behaviour, such as deliberately denying agreed-upon compensation, thus leading to disputes among clients and the cloud server [22]. These incur high costs to the cloud server to monitor such information and may make misjudgments, increasing the difficulty of achieving a sustainable FL system and fair incentive mechanism [15]. In this article, we aim to address the following research question (RQ) in FL.
@@
RQ: "How can we design an incentive mechanism to attract rational clients with highest-quality contributions to improve the FL learning and training efficiency, and mitigate information asymmetry between clients and the cloud server, given the crucial role of CLPs in determining the final FL model performance?"
@@
To address these aforementioned challenges, we propose a CLP-aware incentive mechanism, named Right Reward Right Time (R3T), which leverages contract theory to effectively handle the information asymmetry issues [21]. In addition, R3T integrates with blockchain smart contract to handle the issues of centralized incentive governance and disputes caused by relying on the cloud server [23, 24, 25]. Our approach is supported with theoretical modeling and analysis, together with intensive experiments. Specifically, a set of tailored contract items is designed by the cloud server for clients with different training capabilities per training round. Realizing the importance of CLPs, the cloud server optimally prioritizes and provides corresponding generous rewards for clients who joined early and contributed efforts in the early training phase under complete and incomplete information cases. By providing the right reward at the right time, R3T enables highest-quality clients to join early and contribute efforts for learning-based tasks given by the existence of CLPs in FL thanks to open, transparent, and fair contract items executed on blockchain smart contracts.
1.2 Contributions
@@
    CLP-aware incentive mechanism design: We propose a time-aware contract theory-based model, called R3T, to characterize the cloud utility as a function of rational client’s time and system capabilities, effort, joining time, and reward. Then, we develop a CLP-aware incentive solution and algorithm to jointly optimize client strategies and cloud utility while adaptively attracting high-quality rational clients in the early training stages to enhance FL training and learning efficiency. The proposed incentive mechanism provides the right reward at the right time and handles issues of information asymmetries between clients and the cloud server.
@@
    Performance evaluation: We carry out simulations to verify the effectiveness of our R3T in information symmetry and information asymmetry cases under individual rationality, incentive compatibility, and budget feasibility constraints. R3T demonstrates feasibility and superior efficiency compared with conventional incentive mechanism benchmarks, including linear pricing and contract theory-based methods.
@@
    Proof of concept: We develop a proof of concept of R3T and demonstrate the feasibility and efficiency of the proposed mechanism and system design. Significantly, we show that R3T can boost the convergence time by up to 300% while achieving competitive final FL model accuracies on standardized benchmark datasets, CIFAR-10 and Fashion-MNIST, compared with the state-of-the-art incentive mechanism benchmarks.
1.3 Related Work
Critical learning periods in artificial neural networks. Achille et al. in [5] empirically investigated the correlation between critical periods in the memorization phase and the training outcome in deep neural networks. Their findings demonstrated that analogous to animals, sensory deficits in CLPs cause lasting impairments to the information and connectivity of the network, as quantified by an approximation of the Fisher Information, and adversely affect learning outcomes regardless of subsequent additional training. Similarly, Kleinman et al. in [26] analytically found that the depth of the deep neural network model and data distribution structure are fundamental sources of CLPs in deep neural networks, not only being explained by ascribing to biological ageing processes observed in animals. A similar pattern was found in the distributed training manner, where Yan et al. in [3, 11] showed that FL exhibits CLPs, which play a vital role in determining the final training performance, given the complex of statistical and system heterogeneity. To further explain the phenomenon, metrics such as federated Fisher Information matrix and CLP-aware federated norm vector are proposed to trace and detect CLPs. However, in FL settings, these CLP-given studies are based on the assumption that clients act without cost compensation, rationality, self-interest, and strategic behavior, which may not accurately reflect real-world conditions. Without proper CLP-aware incentives, it may lead to the lack of high-quality contributions in CLPs that have been shown to be permanent and irrecoverable impairment on the global model performance, despite subsequent efforts.
Incentive mechanism design. By incorporating incentive mechanisms into FL systems, engagement and cooperation are promoted where clients are rational, resulting in enhanced learning and training efficiency [27]. Realizing the importance of incentive mechanisms, there are several efforts to develop game-theoretic incentive techniques in FL (e.g., [28, 29, 30, 31, 32, 33, 34]). Authors in [28, 29, 30] designed leader-follower-based incentive mechanisms to encourage clients’ contribution, where the cloud server sets the total reward or price per contribution unit, and then FL clients determine contribution strategies (e.g., data quantity or local performance) to maximize their utilities. However, the information asymmetry condition was not considered in these studies. To handle this issue, utilizing the self-revelation principle of contract theory [21], Sun et al. in [31] proposed an incentive mechanism using a single client’s property (i.e., privacy leakage costs) to create different contract items under both complete and incomplete information cases to compensate for client’s privacy leakage while ensuring comparable learning model performance. Under the same information cases, Huang et al. in [33] proposed a data quality-based incentive mechanism, where the client’s data size, data quality, and unit data computing cost are considered, to derive optimal contract items to optimize the cloud server’s costs in the era of artificial intelligence-generated. Tang et al. in [34] proposed a non-cooperative game-based incentive mechanism to model the interactions between clients (i.e., organizations), deriving optimal processing capacity of each client, to address free-rider problems of public goods in FL under constraints of individual rationality and budget balance. To provide further comprehensive incentive solutions by handling the transparency, centralization, and potential disputes among clients and the cloud server, blockchain is fused with FL systems to manage incentive calculation and distribution [29, 35, 30, 36, 37, 34]. For instance, Yuan et al. in [37] utilized Ethereum-based smart contracts to autonomously execute an incentive mechanism, that considers influences from competition and provides fair and non-repudiative compensation and incentive, to improve social welfare. Nevertheless, existing works on incentive mechanisms in FL to provide the right reward at the right time have not been explored yet. None of these works considers the existence of CLPs, which have shown to play a vital role in determining final FL model performance, and these studies treat every training round as equally important. Besides, information asymmetries among clients and the cloud server, and their impacts on CLPs’s advantages and the final model performance have not been analyzed yet.
2 System model
2.1 System Overview
@@
Figure 2: R3T system’s learning workflow: (1) prepare contracts, (2) sign contracts & download global model, (3) local training, (4) upload updates, (5) forward & validate updates, (6) receive & aggregate model updates, and (7) settlement.
@@
As depicted in Fig. 2, the R3T system consists of three main subsystems: edge servers forming a blockchain platform, clients, and a cloud server. We denote a set of edge servers as $\mathcal{L}$, where each edge server $l\in\mathcal{L}$ connects with clients to exchange information. $\mathcal{L}$ forms, maintains, and operates a blockchain platform to record and validate information such as client identification (ID), local model updates, global models, joining time, and effort levels. Also, R3T integrates a distributed file storage system such as InterPlanetary File System (IPFS), to achieve privacy preservation, high availability, scalability, and low latency for data storage and retrieval [38]. The set of clients is denoted as $\mathcal{N}=\{1,...,n,...,N\}$, where $N$ represents the number of clients. Clients join and contribute their time, computing, and data resources to jointly train the global model published by the cloud server, in exchange for benefits such as monetary rewards [39]. In each training round, a subset of clients $\mathcal{S}_{n}\subseteq\mathcal{N}$, participates in improving the model performance (e.g., in the training round $t$, the cloud server randomly selects $\mathcal{S}_{n}^{t}$ clients out of $\mathcal{N}$). In addition, each client owns or has a right to use or share its dataset $\mathcal{D}$, and the local loss function of a client is defined as ${F}$. In the system, the cloud server plays roles of initiating the global model parameters ${w}$ and seeking high-quality clients to optimize the model performance. It receives local model updates from clients, conducts model aggregation, and handles payments via a blockchain platform.
2.2 Learning Workflow

### v2

@@ Abstract
Critical learning periods (CLPs) in federated learning (FL) refer to early stages during which low-quality contributions (e.g., sparse training data availability, strategic non-participation) can permanently impair the performance of the global model owned by the model owner (i.e., a cloud server). However, existing incentive mechanisms typically exhibit temporal homogeneity, treating all training rounds as equally important, thereby failing to prioritize and attract high-quality contributions during CLPs. This inefficiency is compounded by information asymmetry due to privacy regulations, where the cloud lacks knowledge of client training capabilities, leading to adverse selection. Thus, in this article, we propose a time-aware contract-theoretic incentive framework, named Right Reward Right Time (R3T), to encourage client involvement, especially during CLPs, to maximize the utility of the cloud server in FL. We formulate a cloud utility function that captures the trade-off between the achieved model performance and rewards allocated for clients’ contributions, explicitly accounting for client heterogeneity in time and system capabilities, effort, and joining time. Then, we devise a CLP-aware incentive mechanism deriving an optimal contract design that satisfies individual rationality, incentive compatibility, and budget feasibility constraints, motivating rational clients to participate early and contribute efforts. By providing the right reward at the right time, our approach can attract the highest-quality contributions during CLPs. Simulation and proof-of-concept studies show that R3T mitigates information asymmetry, increases cloud utility, and yields superior economic efficiency than conventional incentive mechanisms benchmarks. Our proof-of-concept results demonstrate up to a 47.6% reduction in the total number of clients and up to a 300% improvement in convergence time while achieving competitive test accuracy.
I Introduction
I-A Background and Motivations
@@
Federated learning (FL) is a privacy-preserving distributed learning paradigm, where distributed clients (e.g., devices or organizations) collaboratively train a shared artificial intelligence model under the coordination of a central server without disclosing private data [mcmahan2017communication, nguyen2025federated]. In classical FL settings, the importance of all training rounds and the efforts of distributed clients across these rounds are weighted equally. However, this setting is questioned by a recent finding in FL known as critical learning periods (CLPs), which highlight the varying impacts of different training rounds and significantly affect FL learning and training efficiency [yan2022seizing, huang2025pa3fed]. This finding was inspired by earlier research on CLPs in species’ cognitive and learning functions (i.e., humans and animals) and deep learning networks within a centralized setting [kandel2000principles, achille2018critical, jastrzebski2021catastrophic, golatkar2019time]. Particularly, CLPs are early learning phases (c.f., Fig. 1(1(a))) affecting long-term learning capabilities [hensch2005critical]. Empirical studies show that the absence of proper learning during these periods can lead to irreversible deficits. For instance, barn owls exposed to misaligned auditory and visual cues during their critical development periods fail to properly localize spatial locations in adulthood [knudsen1990sensitive]. This phenomenon was also observed in deep neural networks, in which any deficits and low effort (e.g., sparse data availability, late client participation) during this initial learning stage can cause lasting impairment on model performance, regardless of subsequent efforts (cf., Fig. 1(1(b))) [kleinman2023critical, yan2022seizing]. The reason for the phenomenon is still unclear, but possible explanations include the depth of the learning model, the structure of training data distribution, or defects in the implementation and training process [achille2018critical, jastrzebski2021catastrophic, golatkar2019time].
@@
Figure 1: Illustration of: a) CLPs in the biological field [hensch2004critical], and b) how early effort affects FL global model performance [yan2022seizing].
@@
Despite the promising merits of enhancing FL efficiency (e.g., accuracy and convergence rate), CLPs remain under-explored, including a lack of transparent, fair, and timely incentives for participation and contribution elicitation, and information asymmetry among clients and the global model owner (e.g., a cloud server [^1] The cloud server and the cloud are used interchangeably.) during these periods. Existing CLP-considered works in FL, such as CLPs detection, client selection, or CLP-aware aggregation, are based on a common assumption of voluntary engagement of clients in the training process (e.g., [yan2022seizing, yan2023criticalfl, huang2025pa3fed, 11060892]). However, this is impractical because clients are rational, strategic, benefit-driven, and need sufficient rewards to compensate for their efforts (e.g., data, computation resources, and time costs). This leads to a lack of client motivation for contribution, resulting in the unsustainable development of FL systems. Although several works have been proposed to bring incentivization by considering effort-reward approaches to fairly compensate FL clients’ effort (e.g., [kang2019incentive, 10097767, chen2024fdfl, 10323190]), they treat clients’ efforts and their impacts on the global model performance on every training round equally and ignore the temporal factor, whereas efforts in CLPs have more weights and mainly contribute to the overall performance of the global model. By failing to prioritize CLP-driven incentive solutions, current methods inefficiently allocate cloud rewards (e.g., money), thereby reducing cloud utility—defined as the difference between gains from global model performance and compensation costs (which is defined later in Section II).
Furthermore, the design of efficient incentives is hindered by information asymmetry between the FL clients and the cloud server. Specifically, due to privacy restrictions, the cloud lacks prior knowledge of the clients’ private training capabilities, including the data, communication, and computation resources, efforts (e.g., data quantity and distribution), or available joining time. Therefore, assumptions that the cloud server knows precise information [10539084, zhan2020learning, sarikaya2019motivating], such as clients’ computing resources or clients’ data size, may not hold in practice. A compounded challenge overlooked by prior works is that existing contract-theoretic incentive approaches, while effective for handling information asymmetry in FL incentives (e.g., [sun2021pain]), fail to incorporate the temporal heterogeneity of CLPs, where early deficits cause irreversible model impairments regardless of later efforts [yan2022seizing, huang2025pa3fed]. This leads to suboptimal contract designs that treat all training rounds equally, inefficiently allocating incentives without prioritizing high-quality contributions [^2] We refer to high-quality contributions to high training efforts, such as a huge amount of closer-distribution-to-the-learning-task data and large computation resources, and they are available in CLPs. during CLPs, resulting in diluted contributions and permanent performance losses under asymmetry about clients’ training capabilities.
There is also uncertainty regarding whether clients will exert the promised effort post-contract [bolton2004contract]. Even with contractually specified incentives, clients may engage in moral-hazard behaviors (e.g., low-effort training or strategic dropouts), and the cloud may be accused of delaying or reneging on payments, which can lead to disputes. Addressing these issues requires costly monitoring and verification and remains error-prone, increasing the difficulty of sustaining FL and implementing fair incentive mechanisms [10097767]. In this article, we aim to address the following research question (RQ).
@@
RQ: How can we design an incentive mechanism to attract rational clients with highest-quality contributions to improve FL learning and training efficiency, and mitigate information asymmetry between clients and the cloud, given the crucial role of CLPs in determining the final FL model performance?
@@
To address the aforementioned challenges, we propose a CLP-aware incentive contract-theoretic framework, named Right Reward Right Time (R3T). Leveraging contract theory, R3T effectively mitigates the information asymmetry issues by designing a set of temporally integrated contract items for the cloud server [bolton2004contract]. The optimal contract design incentivizes clients to self-reveal their training capabilities, join early, and contribute efforts by optimally prioritizing and providing corresponding generous rewards in the early training phase under complete and incomplete information cases, while satisfying individual rationality, incentive compatibility, and budget feasibility constraints. By providing the right reward at the right time, R3T enables the highest-quality clients to join early and contribute efforts during CLPs. R3T also integrates with blockchain smart contracts to handle the issues of centralized incentive governance and disputes caused by relying on the cloud server [wang2022infedge, yu2023ironforge, 3718082].
I-B Related Work
Critical learning periods in neural networks. Achille et al. in [achille2018critical] empirically investigated the correlation between critical periods in the memorization phase and the training outcome in deep neural networks. Their findings demonstrated that, analogous to animals, sensory deficits in CLPs cause lasting impairments to the information and connectivity of the network, as quantified by an approximation of the Fisher Information, and adversely affect learning outcomes regardless of subsequent additional training. Similarly, Kleinman et al. in [kleinman2024critical] analytically found that the depth of the deep neural network model and data distribution structure are fundamental sources of CLPs in deep networks, not only being explained by ascribing to biological aging processes observed in animals. A similar pattern was found in the distributed training manner, where Yan et al. in [yan2022seizing, yan2023criticalfl] showed that FL exhibits CLPs, which play a vital role in determining the final training performance, given the complexity of statistical and system heterogeneity. To further explain the phenomenon, metrics such as federated Fisher Information matrix and CLP-aware federated norm vector are proposed to trace and detect CLPs.
However, these CLP-aware FL studies assume voluntary participation and overlook costs and benefit-driven and strategic behavior, which do not reflect real-world conditions. Without proper CLP-target incentives, it may lead to the lack of high-quality contributions in CLPs that have been shown to be permanent and irrecoverable impairment on the global model performance, despite subsequent efforts.
Incentive mechanisms. By incorporating incentive mechanisms into FL systems, engagement and cooperation are promoted where clients are rational, resulting in enhanced efficiency [anonymous2026incentives]. Realizing the importance of incentive mechanisms, there are several efforts to develop game-theoretic incentive techniques in FL (e.g., [10264195, 9497718, han2024dynamic, sun2021pain, ding2023joint, 10557146, 10418498]). Authors in [10264195, 9497718, han2024dynamic] designed leader-follower-based incentive mechanisms to encourage clients’ contribution, where the cloud server sets the total reward or price per contribution unit, and then FL clients determine contribution strategies (e.g., data quantity or local performance) to maximize their utilities. However, the information asymmetry condition was not considered in these studies. To handle this issue, utilizing the self-revelation principle of contract theory [bolton2004contract], Sun et al. in [sun2021pain] proposed an incentive mechanism using a single client’s property (i.e., privacy leakage costs) to create different contract items under both complete and incomplete information cases to compensate for client’s privacy leakage while ensuring comparable learning model performance. Under the same information cases, Huang et al. in [10557146] proposed a data quality-based incentive mechanism, where the client’s data size, data quality, and unit data computing cost are considered, to derive optimal contract items to optimize the cloud server’s costs in the era of artificial intelligence-generated. Tang et al. in [10418498] proposed a non-cooperative game-based incentive mechanism to model the interactions between clients (i.e., organizations), deriving optimal processing capacity of each client, to address free-rider problems of public goods in FL under constraints of individual rationality and budget balance.
To provide further comprehensive incentive solutions by handling the transparency, centralization, and potential disputes among clients and the cloud server, blockchain is fused with FL systems to manage incentive calculation and distribution [wang2023incentive, han2024dynamic, yuan2023tradefl, 10418498]. For instance, Yuan et al. in [yuan2023tradefl] utilized Ethereum-based smart contracts to autonomously execute an incentive mechanism, that considers influences from competition and provides fair and non-repudiative compensation and incentive, to improve social welfare.
Nevertheless, existing works on incentive mechanisms in FL to provide the right reward at the right time have not been explored yet. None of these works consider the existence of CLPs, which have been shown to play a vital role in determining final FL model performance, and these studies treat every training round as equally important. Besides, information asymmetries among clients and the cloud server, and their impacts on CLPs’ advantages and the final model performance have not been analyzed yet.
I-C Contributions
@@
    Temporal-integrated contract-theoretic incentive framework: We develop a time-aware contract-theoretic framework, named R3T, which is the first to explicitly integrate CLPs into FL incentive design. R3T characterizes cloud utility as a function of client time and system capability, effort, joining time, and reward decisions.
@@
    CLP-aware incentive mechanism design: We develop a CLP-aware incentive mechanism that analytically derives an optimal contract to jointly optimize client participation strategies and cloud utility while adaptively attracting high-quality rational clients in the early training stages to enhance FL training and learning efficiency. The proposed incentive mechanism provides the right reward at the right time and handles issues of information asymmetries between clients and the cloud.
@@
    Performance evaluation: We conduct simulations to verify the effectiveness of our R3T in information symmetry and information asymmetry cases under individual rationality, incentive compatibility, and budget feasibility constraints. R3T demonstrates feasibility and superior efficiency compared with conventional incentive mechanism benchmarks, including linear pricing and conventional contract theory-based methods.
[…]
@@
    Proof of concept: We implement a proof of concept of R3T and demonstrate the feasibility and efficiency of the proposed mechanism and system design. Significantly, we show that R3T can boost the convergence time by up to 300% while achieving competitive final FL model accuracies on standardized benchmark datasets, CIFAR-10 and Fashion-MNIST, compared with the state-of-the-art incentive mechanism benchmarks.
II System model
II-A System Overview
@@
Figure 2: R3T’s one-round learning workflow: (1) prepare contracts, (2) sign contracts & download global model, (3) local training, (4) upload updates, (5) forward & validate updates, (6) receive & aggregate model updates, and (7) settlement.
@@
As depicted in Fig. 2, the R3T framework consists of three main components: edge servers forming a blockchain platform, clients, and a cloud server. We denote a set of edge servers as $\mathcal{L}$, where each edge server $l\in\mathcal{L}$ connects with clients to exchange information. $\mathcal{L}$ forms, maintains, and operates a blockchain platform to record and validate information such as client identification (ID), local model updates, global models, joining time, and effort levels. Also, R3T integrates a distributed file storage system such as InterPlanetary File System (IPFS), to achieve privacy preservation, high availability, scalability, and low latency for data storage and retrieval [jadav2022blockchain]. The set of clients is denoted as $\mathcal{N}=\{1,...,n,...,N\}$, where $N$ represents the number of clients. Clients join and contribute their time, computing, and data resources to jointly train the global model published by the cloud server, in exchange for benefits such as monetary rewards. In each training round, a subset of clients $\mathcal{S}_{n}\subseteq\mathcal{N}$ participates in improving the model performance (e.g., in the training round $t$, the cloud server randomly selects $\mathcal{S}_{n}^{t}$ clients out of $\mathcal{N}$). Besides, clients have the right to use or share their dataset $\mathcal{D}$, and the local loss function of a client is defined as F{F}. In R3T, the cloud server initiates global model parameters w{w} and seeks high-quality clients to optimize the global model performance. It receives local model updates from clients, conducts model aggregation, and handles payments via a blockchain platform.
II-B Learning Workflow